# Bibliografia.bib

### v1

@@ 4. The rationality and semi-rationality of nilpotent groups
    $\chi^{j}(t^{i}(x_{0},x_{1}\cdots x_{p-1}))=\begin{cases}\chi(x_{0})\chi(x_{1})\cdots\chi(x_{p-1}),&\text{if }i=0;\\ \zeta_{p}^{ij}\chi(x_{0}x_{i}x_{2i}\cdots x_{(p-1)i}),&\text{if }i\neq 0.\end{cases}$
@@ 4.1. Rationality
Then $K$ is the compositum of the fields $K_{p}$ with $p\in\pi(n)$. Therefore, if $G_{p}$ is a $p$-group of exponent $n_{p}$ with ${\mathbb{Q}}(G_{p})=K_{p}$ for each $p\in\pi(n)$, then the group $G=\prod_{p\in\pi(n)}G_{p}$ is nilpotent of exponent $n$ with $K={\mathbb{Q}}(G)$. This reduces the proof to the case where $n=p^{e}$ for some prime $p$ and a positive integer $e$. Then $m=p^{t}$ for some $t\leq e$. If $p$ is odd then $K={\mathbb{Q}}_{p^{t}}$, and, by Corollary 4.2, the group $G=C_{p^{t}}\underbrace{\wr C_{p}\wr\dots\wr C_{p}}_{e-t\text{ times}}$ has exponent $p^{e}$ and ${\mathbb{Q}}(G)=K$, as desired.
@@ 4.2. Semi-rationality
In this subsection we classify, for every positive integer $n$, the cosets of $\mathcal{U}_{n}$ that occurs as the semi-rationality of an USR nilpotent group of exponent $n$. We begin with a result which shows that, while 2.1 demonstrate that no converse of the implications in Figure 1 holds for $2$-groups, for $3$-groups all the properties displayed, except that of being rational, are equivalent.
@@
(8) implies (1) and (2). Suppose that $4\in\mathcal{R}_{G}$. If $G=1$, then it is rational and hence it is both inverse semi-rational and quasi-rational. Otherwise $-1\not\in\mathcal{R}_{G}$ and, as $[\mathcal{U}_{G}:\left\langle 4\right\rangle]=2$, it follows that $\mathcal{R}_{G}=\left\langle 4\right\rangle$. Thus $G$ is quasi-rational and $\mathcal{S}_{G}=\mathcal{U}_{G}\setminus\left\langle 4\right\rangle$, by Proposition 3.13. Hence $-1\in\mathcal{S}_{G}$, i.e. $G$ is inverse semi-rational. Moreover, by equation (2.1), $\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{Q}}(G))=\left\langle\sigma_{4}\right\rangle=\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{Q}}(\zeta_{3}))$ and hence ${\mathbb{Q}}(G)={\mathbb{Q}}(\zeta_{3})$. This concludes the proof.
@@
Now we prove that (2) implies (3). Suppose that $G$ is USR. Then $C_{p}$ is USR, by Remark 3.14, and hence $p$ is $2$ or $3$. By means of contradiction, suppose that $X$ is not quasi-rational. Then there are two irreducible characters $\chi_{1}$ and $\chi_{2}$ of $X$ such that ${\mathbb{Q}}(\chi_{1})$ and ${\mathbb{Q}}(\chi_{2})$ are different quadratic extensions of ${\mathbb{Q}}$, so that there are $x_{1},x_{2}\in X$ with $\chi_{2}(x_{1})\not\in{\mathbb{Q}}(\chi_{1})$ and $\chi_{1}(x_{2})\not\in{\mathbb{Q}}(\chi_{2})$. Consider the character $\chi=(\chi_{1}\chi_{2})^{G}$ of $G$. Observe that this is a character of the first type in Theorem 4.1 because $\chi=(\chi_{1}\chi_{2}\chi_{0}\dots\chi_{0})^{G}$ for $\chi_{0}$ the trivial character of $X$. Then $\chi$ is an irreducible character of $G$. Then $\chi(x_{2},1,1,\dots,1)=\chi_{1}(x_{2})$ and $\chi(1,x_{1},1,\dots,1)=\chi_{2}(x_{1})$. Therefore ${\mathbb{Q}}(\chi_{1}(x_{2}),\chi_{2}(x_{1}))\subseteq{\mathbb{Q}}(\chi)$ and hence $[{\mathbb{Q}}(\chi):{\mathbb{Q}}]\geq[{\mathbb{Q}}(\chi_{1}(x_{2}),\chi_{2}(x_{1})):{\mathbb{Q}}]=4$. Therefore, $G$ is not quadratic rational, in contradiction with Corollary 3.12. Thus $X$ is quasi-rational. If $p=2$, then condition (3a) holds. Now suppose that $p=3$. Then for the irreducible character $\chi^{1}$ of $G$ of the second type in Theorem 4.1, we get $\chi^{1}(t)=\zeta_{3}$ so that ${\mathbb{Q}}(\chi^{1})={\mathbb{Q}}(\zeta_{3})$. Then $\chi(x)=\chi^{1}(x,1,\dots,1)\in{\mathbb{Q}}(\zeta_{3})$ for every $x\in X$. This shows that ${\mathbb{Q}}(X)\subseteq{\mathbb{Q}}(\zeta_{3})$, i.e. condition (3b) holds.

### v2

@@ 4. The rationality and semi-rationality of nilpotent groups
    $\chi^{j}(t^{i}(x_{0},x_{1},\dots,x_{p-1}))=\begin{cases}\chi(x_{0})\chi(x_{1})\cdots\chi(x_{p-1}),&\text{if }i=0;\\ \zeta_{p}^{ij}\chi(x_{0}x_{i}x_{2i}\cdots x_{(p-1)i}),&\text{if }i\neq 0.\end{cases}$
@@ 4.1. Rationality
Then $K$ is the compositum of the fields $K_{p}$ with $p\in\pi(n)$. Therefore, if $G_{p}$ is a $p$-group of exponent $n_{p}$ with ${\mathbb{Q}}(G_{p})=K_{p}$ for each $p\in\pi(n)$, then the group $G=\prod_{p\in\pi(n)}G_{p}$ is nilpotent of exponent $n$ with $K={\mathbb{Q}}(G)$. This reduces the proof to the case where $n=p^{e}$ for some prime $p$ and a positive integer $e$. Then $m=p^{t}$ for some $1\leq t\leq e$. If $p$ is odd then $K={\mathbb{Q}}_{p^{t}}$, and, by Corollary 4.2, the group $G=C_{p^{t}}\underbrace{\wr C_{p}\wr\dots\wr C_{p}}_{e-t\text{ times}}$ has exponent $p^{e}$ and ${\mathbb{Q}}(G)=K$, as desired.
@@ 4.2. Semi-rationality
In this subsection we classify, for every positive integer $n$, the cosets of $\mathcal{U}_{n}$ that occur as the semi-rationality of an USR nilpotent group of exponent $n$. We begin with a result which shows that, while 2.1 demonstrate that no converse of the implications in Figure 1 holds for $2$-groups, for $3$-groups all the properties displayed, except that of being rational, are equivalent.
@@
(8) implies (1) and (2). Suppose that $4\in\mathcal{R}_{G}$. If $G=1$, then it is rational and hence it is both inverse semi-rational and quadratic. Otherwise $-1\not\in\mathcal{R}_{G}$ and, as $[\mathcal{U}_{G}:\left\langle 4\right\rangle]=2$, it follows that $\mathcal{R}_{G}=\left\langle 4\right\rangle$. Thus $G$ is quadratic and $\mathcal{S}_{G}=\mathcal{U}_{G}\setminus\left\langle 4\right\rangle$, by Proposition 3.13. Hence $-1\in\mathcal{S}_{G}$, i.e. $G$ is inverse semi-rational. Moreover, by equation (2.1), $\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{Q}}(G))=\left\langle\sigma_{4}\right\rangle=\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{Q}}(\zeta_{3}))$ and hence ${\mathbb{Q}}(G)={\mathbb{Q}}(\zeta_{3})$. This concludes the proof.
@@
Now we prove that (2) implies (3). Suppose that $G$ is USR. Then $C_{p}$ is USR, by Remark 3.14, and hence $p$ is $2$ or $3$. By means of contradiction, suppose that $X$ is not quadratic. Then there exist $\chi_{1},\chi_{2}\in\operatorname{Irr}(X)$ such that ${\mathbb{Q}}(\chi_{1})$ and ${\mathbb{Q}}(\chi_{2})$ are different quadratic extensions of ${\mathbb{Q}}$, so that there are $x_{1},x_{2}\in X$ with $\chi_{2}(x_{1})\not\in{\mathbb{Q}}(\chi_{1})$ and $\chi_{1}(x_{2})\not\in{\mathbb{Q}}(\chi_{2})$. Consider the character $\chi=(\chi_{1}\chi_{2})^{G}$ of $G$. Observe that this is a character of the first type in Theorem 4.1 because $\chi=(\chi_{1}\chi_{2}\chi_{0}\dots\chi_{0})^{G}$ for $\chi_{0}$ the trivial character of $X$. Then $\chi\in\operatorname{Irr}(G)$. Then $\chi(x_{2},1,1,\dots,1)=\chi_{1}(x_{2})$ and $\chi(1,x_{1},1,\dots,1)=\chi_{2}(x_{1})$. Therefore ${\mathbb{Q}}(\chi_{1}(x_{2}),\chi_{2}(x_{1}))\subseteq{\mathbb{Q}}(\chi)$ and hence $[{\mathbb{Q}}(\chi):{\mathbb{Q}}]\geq[{\mathbb{Q}}(\chi_{1}(x_{2}),\chi_{2}(x_{1})):{\mathbb{Q}}]=4$. Therefore, $G$ is not character quadratic, in contradiction with Corollary 3.12. Thus $X$ is quadratic. If $p=2$, then condition (3a) holds. Now suppose that $p=3$. Then for any irreducible character $\chi^{1}$ of $G$ of the second type in Theorem 4.1, we get $\chi^{1}(t)=\zeta_{3}$ so that ${\mathbb{Q}}(\chi^{1})={\mathbb{Q}}(\zeta_{3})$. Then $\chi(x)=\chi^{1}(x,1,\dots,1)\in{\mathbb{Q}}(\zeta_{3})$ for every $x\in X$. This shows that ${\mathbb{Q}}(X)\subseteq{\mathbb{Q}}(\zeta_{3})$, i.e. condition (3b) holds.
@@
Acknowledgment. The authors thank the reviewer for the careful reading and insightful comments.